A bag contains 19 balls numbered 1 through 10. What is the probability that a randomly selected ball has an even number?
Answer by stanbon(75887)

A bag contains 19 balls numbered 1 through 10. What is the probability that a randomly selected ball has an even number?
----------
# of even numbers is 5: 2,4,6,8,10
# of numbers is 10
Prob(picking an even number)=5/10=1/2
Cheers,
Stan H.